OpenStudy (anonymous):

find the LCD for 3/7+1/3. explain

OpenStudy (anonymous):

i dont really know what can be divisible for 7 and 3 besides 1

OpenStudy (anonymous):

?

geerky42 (geerky42):

LCD is basically like LCM of denominators.

geerky42 (geerky42):

So what is LCM of 3 and 7?

OpenStudy (anonymous):

thats alot of reading...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

idrk cause nothing is really divisible

geerky42 (geerky42):

LCM = least common multiple

OpenStudy (anonymous):

i know that but i dont get how to find that out for 7 and 3

geerky42 (geerky42):

What is prime factorization of 3? What about 7?

OpenStudy (anonymous):

nothing can go into 3 or 7 besides 1

geerky42 (geerky42):

Right, so 3 and 7 are both prime, right?

OpenStudy (anonymous):

yeah

geerky42 (geerky42):

So to find LCM of primes number, you just multiple them together.

geerky42 (geerky42):

So \(LCM(3,7)=3\times7\)

OpenStudy (anonymous):

so it would be 3/21?

geerky42 (geerky42):

\(3\times7 = 21\). That tells us that denominator need to be 21, right? In \(\dfrac{3}{7}\), we need denominator to be 21. What to do?

OpenStudy (anonymous):

cause it has to be a number, euh 3x7?

geerky42 (geerky42):

well, to get from \(7\) to \(21\), we multiply it by \(3\), right? But because we did something in denominator, we must do same to numerator. So we have \(\dfrac{3}{7}\rightarrow\dfrac{3\times3}{7\times3} = \dfrac{9}{21}\). Makes sense? Can you do that for 1/3 ?

OpenStudy (anonymous):

hmmm 1/3? wouldnt it be the same?

geerky42 (geerky42):

Well, to add fraction together, we need denominator to be same, so we need to rewrite 1/3.

OpenStudy (anonymous):

hmm rewrite 1/3. that sounds confusing

geerky42 (geerky42):

rewrite 1/3 means write it again, so it is same thing, but look different.

geerky42 (geerky42):

Like for 1/2, we can rewrite it as 2/4, or 4/8

OpenStudy (anonymous):

ohh let me try

geerky42 (geerky42):

Alright, denominator just need to be 21.

OpenStudy (anonymous):

hmm...

geerky42 (geerky42):

So we need \(\dfrac{1}{3}~\rightarrow~\dfrac{?}{21}\)

OpenStudy (anonymous):

7/21?

geerky42 (geerky42):

What to do with denominator and numerator?

geerky42 (geerky42):

Yes

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

oh so its 7/21?

geerky42 (geerky42):

So we now have \(\dfrac{3}{7}+\dfrac{1}{3}\rightarrow\dfrac{9}{21}+\dfrac{7}{21}\). Now we can add fractions.

OpenStudy (anonymous):

so 9/21 + 7/21?

geerky42 (geerky42):

yeah

OpenStudy (anonymous):

which is 16/21?

OpenStudy (gorv):

yeah that is right

OpenStudy (gorv):

greeky42 guided u well

OpenStudy (anonymous):

lol yeah @geerky42 did

geerky42 (geerky42):

lol thanks, so do you understand what we did so far?

OpenStudy (anonymous):

grrr it has to be a single digit number, but yeah i kinda understand it

OpenStudy (anonymous):

how do i get one number?

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

but i have a few more

OpenStudy (anonymous):

hold on